OpenStudy (anonymous):

solve the equation and write complex solutions in standard form. x^2+16x+74=0

OpenStudy (campbell_st):

use the quadratic formula \[x = (-b \pm \sqrt{b^2 - 4ac})/2a\] in your question a = 1, b = 16, c = 74 the value of \[b^2 -4ac = 16^2 - 4\times1\times74 = -40\] this can be written as \[-40 = 40i^2\] you'll need to simplify the surd